A 10% chlorine mixture is to be mixed to 25% of chlorine mixture to obtain 30 gallon of 20 % how many gallons of each can be used ?
Answer by dkppathak(439)

A 10% chlorine mixture is to be mixed to 25% of chlorine mixture to obtain 30 gallon of 20 % how many gallons of each can be used ?
Answer
let 10% chlorine mixture is X liters
let 25% chlorine mixture is y liters
as per given conditions
0.1 X + 0.25 Y = 6 (1)
X + Y =30 (2) by multiplying (2) by 0.1
0.1 X + 0.25 Y = 6 (3)
0.1 x + 0.1 y = 3 (4)
by subtracting
0.15 y = 3
y =3/0.15 =300/15 =20 gallon
X + Y = 30
X + 20 = 30
X =10 gallon
Answer 10% chlorine mixture 10 gallon
25% chlorine mixture should be 20 gallon